# Problem Solving Through Human-AI Preference-Based Cooperation

### v2

@@ 1 Introduction
To effectively support human-AI cooperation, we draw inspiration from how humans collectively address complex problems to devise solutions: Humans often solve complex problems by iteratively co-constructing a solution step by step, revising and refining draft solutions while transitioning between different levels of abstraction, and exchanging information about preferences and potential improvements in natural language (Damşa, 2013). Our position is that this type of human-human cooperation is a promising template for the cooperation of humans and AI agents. On this basis, we propose Human-AI Co-Construction (HAI-Co2), a novel framework for human-AI cooperative problem-solving that builds on preference-based learning and search methodology and relies on natural language to facilitate interaction. The problem-solving process is conceived as a process of systematic search in a construction space $\mathcal{X}$ of candidate solutions, i.e., as a co-constructive process, in which candidate solutions are modified step by step until a sufficiently good solution has been found.
@@
What are co-construction problems? The broad class of problems that we seek to address in this paper is primarily centered around domain-specific expert applications where the task is to construct a solution that meets specific requirements, e.g., a computer program in software engineering or a machine learning pipeline in automated machine learning (AutoML). We choose the term co-construction because of the necessity for the AI agent to work closely with human experts. First, current AI systems’ capabilities are limited and cannot fully replace human expertise in these tasks, e.g., due to insufficient knowledge and reasoning capabilities, lack of trustworthiness and bias. Second, some expert tasks (e.g., designing websites and building data analysis and visualization software) require irreplaceable human supervision; solutions to be co-constructed in these problems are defined based on personalized human preference. While the co-constructed solution must fulfill a set of objective correctness criteria, the human expert imposes a broad set of subjective criteria as well. For example, in a software development setup, the final piece of software must be executable, secure, and bug-free; at the same time, every expert developer has their own style – of coding, commenting, modularization, etc. – that should be followed as well. This implies that the AI agent must be able to follow and adapt to the personalized preferences of the human expert (which are not necessarily stable but may evolve in the course of the problem-solving process).
@@
Is the current state of generative AI enough? Multiple prior investigations have laid out the inherent shortcomings of present day generative AI that limit its applicability to the type of problems we are addressing in this paper. In an example case study tailored towards code generation presented in Figure 1, we demonstrate some of the major bottlenecks of GPT-4 Turbo, a recent update [^1] We use the gpt-4-1106-preview instance: https://openai.com/index/new-models-and-developer-products-announced-at-devday/ over the original GPT-4 model (OpenAI, 2024). In Figure 1 (a), GPT-4 Turbo ignores the human expert’s explicit instructions to generate a complete Python code with the required module specifications, echoing Xie et al. (2024)’s observation that current language-based AI agents lack task adherence. After repeated prompting with partial code snippets, the process produces a complete – albeit faulty – code. This limitation becomes even more irreconcilable as more varied and realistic expressions of human preference are taken into account – for the human expert to contribute productively, one must allow preferences expressed via explicit instructions, binary choice, ranking, etc. Current generative AI solutions do not facilitate such multi-modal preference incorporation. Figure 1 (b) shows unreliable debugging attempts. Specifically, the LLM performs unrelated (and faulty) edits to address a bug and even introduces new errors. This demonstrates that existing LLMs fail to handle complex, modular software code (Jiang et al., 2024). The common practice is that the human (as a knowledgable expert who keeps track of overall context) identifies faulty output and repeatedly prompts the model to guide it to the correct generation – this is implicitly adopting a co-construction paradigm. However, Figure 1 (c) shows that current modes of human-AI interaction cannot unleash the full potential of co-construction – direct modification of the co-constructed candidate solution by the human expert does not bear any special significance to the LLM and it treats it as just another context. There is no explicit mechanism for the AI to learn implicit preferences expressed by the human through active participation.
@@
While these examples are focused on code synthesis, there is evidence of similar shortcomings in other domains (Kambhampati et al., 2024; Lecler et al., 2023). Code synthesis in particular – and the experience from day-to-day use of generative AI for solving complex problems in general – point toward co-construction as a naturally evolved problem-solving paradigm where the human expert tries to search for the optimal solution by interacting with the AI. However, the current state of generative AI hinders its role as a reliable partner in successful co-construction. This is because the “one-dimensional” interaction between human and AI typical of how AI agents are used today often fails to steer the co-construction towards a solution that satisfies the user’s constraints.
@@
Our contribution. In this paper, we formalize co-constructice problem solving and thereby aim to address important limitations of current generative AI models. We present HAI-Co2, a framework with three fundamental properties that facilitate human-AI co-construction. First, it introduces multiple levels of abstractions to the candidate solution, providing a seamless interface for the human expert and the AI agent to modify and keep track of the complex, modular co-constructed candidate solution. Second, HAI-Co2 allows multi-modal preference input from the human expert, with natural language as the central mediator to capture information-rich guidance signals, along with other forms of active expert participation such as categorical choice-based preference. Finally, HAI-Co2 introduces a search-based methodology of co-construction where the candidate solution (represented on multiple levels of abstraction) is iteratively revised to maximize the perceived utility modeled from the preference input.
@@
The rest of the paper is organized as follows. Following a discussion of relevant literature in Section 2, we introduce the framework of HAI-Co2 in Section 3 and outline open research challenges implied by this framework in Section 4. Section 5 illustrates HAI-Co2 in a case study, prior to concluding the pape in Section 6.
@@ 2 Related work
We group relevant prior work into five major strands: reinforcement learning from human feedback, RLHF search- and evolution-driven construction, assistance games, LLM agents for complex problem solving and persistent solution spaces for iterative construction. We now discuss how they attempt to address expert-AI co-construction. However, as we will see in Section 4, they fail to comprehensively tackle its challenges.
@@
Assistance Games. Hadfield-Menell et al. (2016) introduce the framework of cooperative inverse reinforcement learning, later called assistance games (Shah et al., 2021; Laidlaw et al., 2024), for cooperative decision-making between an AI agent and a human expert. They model human-AI interaction as a two player game betwen human expert and AI agent. The AI agent aims to maximize the expert’s utility function without explicit knowledge of it. The AI agent learns this function by observing the expert’s actions and receiving feedback on its own actions. Our setting is similar in that AI and human collaborate to construct a solution that only the human expert can evaluate. It differs in that we focus on search-based co-construction of solutions rather than decision-making in a cooperative setting. When viewing solution construction as a sequential decision-making problem, our framework can be seen as a search-based approach to assistance games where the AI agent’s actions are the steps taken to construct the solution. Despite the nice theoretical properties of the assistance games framework (Shah et al., 2021), it has not yet been widely applied to real-world problems, likely due to the complexity of the human-AI interaction and the difficulty of interactively modeling the human expert’s utility function from various sources of feedback. We hope that our framework can provide a more practical approach to human-AI co-construction by focusing on the specific challenges of solution construction and leveraging the extensive prior knowledge of pretrained language models to guide the search process.
@@
LLM agents for complex problem solving. The rapid increase in the capabilities of LLMs has triggered multiple recent efforts to integrate them at the core of autonomous agents that interact with the environment, plan, and act to solve complex problems (Wang et al., 2024). Typical approaches adopt integrating different tool-usage capabilities into LLMs via efficient prompting, often with multimodal capabilities (Chen et al., 2023). A single agent is often insufficient to solve complex problems; thus, multiple agents with different capabilities have to be integrated. Recent efforts in LLM-based multi-agent systems seek to mimic such cooperative problem-solving by role-playing LLMs via in-context examples (Li et al., 2023) or fine-tuning (Juneja et al., 2024). Despite some promising achievements, such agentic ecosystems are inherently limited by the constituent LLMs’ inability to plan and execute tasks (Xie et al., 2024; Kambhampati et al., 2024). These frameworks of LLM-based autonomous agents are largely designed for autonomous problem solving, not for human-AI co-construction as we envision. Recently, arguments in favor of strategically allocating tasks between humans and LLM-based agents to exploit their distinct strengths have been put forward (He et al., 2024), which aligns with our approach of leveraging the strengths of both humans and AI agents in co-construction. An alternative view of our framework is hence an extension of LLM-based multi-agent systems with a human agent as a key component, focusing on the co-construction of solutions.
@@
Persistent Solution Space for Iterative Construction. A fundamental component of our proposed framework is the explicit representation of the construction space for systematic solution search. Such a persistent memory can be useful for LLM agents. Sumers et al. (2024) propose a cognitive architecture for language agents that connects LLMs to internal memory and external environments, grounding them in existing knowledge or external observations. Similarly, Modarressi et al. (2024) introduce a structured memory component that LLM agents can use for storage and retrieval. Although these approaches do not directly address expert-AI co-construction challenges, they relate to our approach by providing agents with persistent memory to store intermediate solutions and relevant information for problem solving.
@@ 3 HAI-Co2: Human-AI co-construction through preference-based search
In this section, we propose HAI-Co2, a framework for cooperative problem solving. Broadly speaking, HAI-Co2 is meant to formalize an interactive problem-solving scenario, in which a human expert seeks to (co-)construct a candidate solution – such as a computer program – with the help of an AI agent. The problem-solving process is conceived as a process of systematic search in a space $\mathcal{X}$ of candidate solutions, i.e., as a (co-)constructive process, in which candidate solutions are modified or extended step by step until a (sufficiently) good solution has been found. Therefore, we also refer to the search space $\mathcal{X}$ as the construction space. The construction space, its hierarchical organization and its topology (or neighborhood structure) are depicted in Figure 2.
@@ 3.1 Construction space and abstraction hierarchy
The construction space will typically be large, most often even (countably) infinite. For example, the construction space may consist of all computer programs in a specific programming language. Spaces of this kind cannot be specified in an explicit way. Instead, they will be defined implicitly and may even be adapted or designed on-the-fly in the course of the problem-solving process. In this regard, the formal representation of candidate solutions is of major importance and will strongly influence the efficacy and efficiency of the problem-solving process. Moreover, it is also clear that the representation of solutions will not be universal but rather specific to the expert domain. For example, a computer program will not be represented in the same way as a machine learning pipeline or data science workflow. It should be noted that we do not make any assumption of completeness for candidate solutions: at any stage of the search, a candidate solution ${\bm{x}}\in{\mathcal{X}}$ can be partial or incomplete (i.e., an incomplete codebase, an incomplete ML pipeline, etc.).
During problem solving, it is often useful to look at (candidate) solutions on multiple levels of abstraction. In many cases, for example, a rough draft of the solution is found in a first phase of the process, and this draft is then worked out in more detail in a second phase. More generally, one can imagine a search process that switches back and forth between different levels of abstraction whenever appropriate. Therefore, we assume the construction space ${\mathcal{X}}$ is equipped with a hierarchy of abstraction levels. Formally, this can be modeled by a sequence ${\mathcal{X}}_{0},{\mathcal{X}}_{1},\ldots,{\mathcal{X}}_{J}$ of spaces, where ${\mathcal{X}}_{j}$ is a refinement of ${\mathcal{X}}_{j-1}$ – or, vice versa, ${\mathcal{X}}_{j-1}$ an abstraction of ${\mathcal{X}}_{j}$.
We describe the abstraction process from $\mathcal{X}_{j}$ to $\mathcal{X}_{j-1}$ as a surjection $f_{j}:{\mathcal{X}}_{j}\rightarrow{\mathcal{X}}_{j-1}$ such that ${\bm{x}}^{\prime}=f_{j}({\bm{x}})\in{\mathcal{X}}_{j-1}$; that is, ${\bm{x}}^{\prime}$ is the abstraction of ${\bm{x}}$ on the abstraction level modeled by ${\mathcal{X}}_{j-1}$. We denote by $f^{(-1)}_{j}({\bm{x}}^{\prime})=\{{\bm{x}}\in{\mathcal{X}}_{j}\,|\,f_{j}({\bm{x}})={\bm{x}}^{\prime}\}$ the set of all refinements of ${\bm{x}}^{\prime}\in{\mathcal{X}}_{j-1}$ on abstraction level ${\mathcal{X}}_{j}$. Note that refinements are not unique, which is why a transition from ${\mathcal{X}}_{j-1}$ to ${\mathcal{X}}_{j}$ may come with a certain arbitrariness. In our case study, we work with three levels of abstraction, considering a Python program as a refinement of a UML diagram, which in turn is a refinement of a natural language description.

### v3

@@ 1 Introduction
To effectively support human-AI cooperation, we draw inspiration from how humans collectively address complex problems to devise solutions: Humans often solve complex problems by iteratively co-constructing a solution step by step, revising and refining draft solutions while transitioning between different levels of abstraction, and exchanging information about preferences and potential improvements in natural language (Damşa, 2013). Our position is that this type of human-human cooperation is a promising template for the cooperation of humans and AI agents. On this basis, we propose Human-AI Co-Construction (HAI-Co2), a novel framework for human-AI cooperative problem solving that builds on preference-based learning and search methodology and relies on natural language to facilitate interaction. The problem solving process is conceived as a process of systematic search in a construction space $\mathcal{X}$ of candidate solutions, i.e., as a co-constructive process, in which candidate solutions are modified step by step until a sufficiently good solution has been found.
@@
What are co-construction problems? The broad class of problems that we seek to address in this paper are primarily centered around domain-specific expert applications where the task is to construct a solution that meets specific requirements, e.g., a computer program in software engineering or a machine learning pipeline in automated machine learning (AutoML). We choose the term co-construction because of the necessity for the AI agent to work closely with human experts.111 Not all problems require co-construction; some have unique solutions independent of human feedback. As AI models become more capable, they will be able to solve such problems, which currently require co-construction due to limited AI capability, without help from the human expert since they can better utilize prior knowledge and task specifications. However, in many cases, task descriptions remain inherently incomplete due to either ambiguous specifications or evolving expert preferences, necessitating co-construction with human input regardless of model capability. First, capabilities of current AI systems are limited and cannot fully replace human expertise in these tasks, e.g., due to insufficient knowledge and reasoning capabilities, lack of trustworthiness, and bias. Scaling generative AI systems, particularly language models, has demonstrated improved performance across a range of tasks, e.g., mathword problems and commonsense reasoning. However, even the most powerful LLMs show a lack of robustness under different semantic perturbations that would not have fooled an otherwise robust reasoning machine (Li et al., 2024). While it is unwise to rule out future improvements, their current limitation certainly calls for interventions beyond scaling. Second, some expert tasks (e.g., designing websites and building data analysis and visualization software) require irreplaceable human supervision; solutions to be constructed in these problems are defined based on personalized human preferences. While the co-constructed solution must fulfill a set of objective correctness criteria that can be verified using symbolic tools without any human intervention, the human expert imposes a broad set of subjective criteria as well. For example, in a software development setup, the final piece of software must be executable, secure, and bug-free; at the same time, every expert developer has their own style – of coding, commenting, modularization, etc. – that should be followed as well. Similarly, an ML pipeline developer may want a suboptimal classifier to save computational cost and go for optimality in a later stage of development. This implies that the AI agent must be able to follow and adapt to the personalized preferences of the human expert (which often are not stable, but evolve in the course of the problem solving process).
@@
Is the current state of generative AI enough? Multiple prior investigations have laid out the inherent shortcomings of present-day generative AI that limit its applicability to the type of problems we are addressing in this paper. In an example case study tailored towards code generation presented in Figure 1, we demonstrate some of the major bottlenecks of GPT-4 Turbo, a recent update [^2] We use the gpt-4-1106-preview instance: https://openai.com/index/new-models-and-developer-products-announced-at-devday/ over the original GPT-4 model (OpenAI, 2024). In Figure 1 (a), GPT-4 Turbo ignores the human expert’s explicit instructions to generate a complete Python code with the required module specifications, echoing Xie et al. (2024)’s observation that current language-based AI agents lack task adherence. After repeated prompting with partial code snippets, the process produces a complete – albeit faulty – code. This limitation becomes even more irreconcilable as more varied and realistic expressions of human preferences are taken into account – for the human expert to contribute productively, one must allow preferences expressed via explicit instructions, binary choice, ranking, etc. Current generative AI solutions do not facilitate such multi-modal preference incorporation. Figure 1 (b) shows unreliable debugging attempts. Specifically, the LLM performs unrelated (and faulty) edits to address a bug and even introduces new errors. This demonstrates that existing LLMs fail to handle complex, modular software code (Jiang et al., 2024). The common practice is that the human (as a knowledgeable expert who keeps track of overall context) identifies faulty output and repeatedly prompts the model to guide it to the correct generation – this is implicitly adopting a co-construction paradigm. However, Figure 1 (c) shows that current modes of human-AI interaction cannot unleash the full potential of co-construction – direct modification of the co-constructed candidate solution by the human expert does not bear any special significance to the LLM, and it treats it as just another context. There is no explicit mechanism for the AI to learn implicit preferences expressed by the human through active participation.
@@
While these examples are focused on code synthesis, there is evidence of similar shortcomings in other domains (Kambhampati et al., 2024; Lecler et al., 2023; Almarie et al., 2023). Carroll et al. (2019), for example, demonstrate the necessity of incorporating explicit “human-awareness” in a version of the collaborative game Overcooked, providing evidence that agents fail to coordinate with human subjects without such awareness. Code synthesis in particular – and the experience from day-to-day use of generative AI for solving complex problems in general – points toward co-construction as a naturally evolved problem solving paradigm where the human expert tries to search for the optimal solution by interacting with the AI. However, the current state of generative AI hinders its role as a reliable partner in successful co-construction. This is because the “one-directional” interaction between human and AI typical of how AI agents are used today often fails to steer the co-construction towards a solution that satisfies the user’s constraints.
@@
Our contribution. In this paper, we formalize co-constructive problem solving and thereby aim to address important limitations of current generative AI models. We present HAI-Co2, a conceptual framework with three fundamental properties that facilitate human-AI co-construction. First, it introduces multiple levels of abstractions to the candidate solution, providing a seamless interface for the human expert and the AI agent to modify and keep track of the complex, modular, co-constructed candidate solution. Second, HAI-Co2 allows multi-modal preference input from the human expert, with natural language as the central mediator to capture information-rich guidance signals, along with other forms of active expert participation, such as categorical choice-based preference. Finally, HAI-Co2 introduces a search-based methodology of co-construction where the candidate solution (represented on multiple levels of abstraction) is iteratively revised to maximize the perceived utility modeled from the preference input. While multiple components of HAI-Co2 have been explored in prior research independently across different domains, we are the first to bring them together under a unified conceptual umbrella and to show that jointly with current generative AI, they have the potential to address multiple major challenges in solving expert domain problems. We showcase the efficacy of HAI-Co2 using an example implementation for the expert domain of software engineering.
@@
The rest of the paper is organized as follows. After a discussion of relevant literature in Section 2, we introduce the framework of HAI-Co2 in Section 3 and outline open research challenges implied by this framework in Section 4. Section 5 illustrates HAI-Co2 in a case study, prior to concluding the paper in Section 6.
@@ 2 Related work
We group relevant prior work into five major strands: reinforcement learning from human feedback (RLHF), assistance games, search- and evolution-driven construction, LLM agents for complex problem solving, and persistent solution spaces for iterative construction. We discuss how these approaches attempt to address expert-AI co-construction. However, as we will see in Section 4, they fail to comprehensively tackle these challenges.
@@
Assistance Games. Another approach rooted in reinforcement learning is assistance games (Shah et al., 2021; Laidlaw et al., 2024), originally introduced as cooperative inverse reinforcement learning (Hadfield-Menell et al., 2016). As a framework for cooperative decision-making between an AI agent and a human, assistance games model human-AI interaction as a cooperative game with the agent and the human as players. The AI agent aims to maximize the human’s utility function without explicit knowledge of it. It learns this function by observing the human’s actions and receiving feedback on its own actions, which may include queries to the human.
The process of co-construction in the sense of HAI-Co2 could in principle be seen as an assistive sequential decision-making problem, too, with the expert’s and the agent’s actions being steps taken to construct the solution. However, in practice, casting co-constructive problem solving as an assistance game does not seem particularly useful to us. Roughly speaking, for HAI-Co2 we consider co-constructive problem solving as a search rather than a reinforcement learning task. While both tasks are related, RL is a very general setting that comes with additional challenges, notably the learning of a model through exploration (in a non-deterministic environment with unknown effects of actions). Indeed, the price that RL pays for its generality is a high conceptual and computational complexity. This problem applies in particular to assistance games, which, despite their nice theoretical properties (Shah et al., 2021), have not yet been widely applied to real-world problems. [^3] Laidlaw et al. (2024) present a first step towards scalably solving assistance games in more complex environments. This is early work, though, and we believe that the conceptual and computational benefits of the search-based formulation justify a more specialized framework for this highly relevant use case. We hope that our framework can provide a more practical approach to human-AI co-construction by focusing on the specific challenges of solution construction with a search process guided by the extensive prior knowledge of pretrained language models.
Learning from natural language interactions. Natural language interaction between the expert and the AI agent is central to HAI-Co2. A popular approach towards facilitating human-AI interactive problem solving involves training the agent to follow instructions in natural language (Branavan et al., 2009; Tellex et al., 2011). This paradigm of learning to follow instructions has found attention in the LLM era as well (Wei et al., 2022). However, directly mapping language-specified goals to actions has limited applicability to novel tasks. Alternatively, learning rewards from natural language interaction to successfully align the AI agent with the human user has been explored (Fu et al., 2019; Sumers et al., 2021) – instead of learning to map language-defined goals to actions directly, they seek to learn the reward function from the language-defined goals that can be generalized to novel tasks. The findings of Sumers et al. (2022) suggest that while instructions typically perform well in low-autonomy settings, high-autonomy regimes favor the reward-learning paradigm. Instead of learning a language-conditioned policy or reward, it is also possible to use conversational cues as rewards themselves (Jaques et al., 2020), which can be combined with the other approaches discussed here. Yet another approach, deployed in OpenAI’s ChatGPT, is to enable the language model to save information about the user and their interactions with the model as natural-language ‘memories’, which may include information about the user’s preferences. [^4] https://openai.com/index/memory-and-new-controls-for-chatgpt/ Most of these approaches do not consider the need to actively elicit and co-construct the expert’s preferences, a key aspect of HAI-Co2. This necessity is supported by Co-Reyes et al. (2019) and Lin et al. (2022), who identify that inferring the correct behavior or reward from a single utterance can become non-trivial given the multidimensionality of language. Querying the human user and estimating their preferences in an interactive setup is also a key component of Peng et al. (2024)’s framework. As a step in this direction, Li et al. (2023a) use active elicitation to strengthen preference understanding – the AI agent is trained to elicit and infer human preferences by actively interacting with the user. This is crucial prior work for an implementation of HAI-Co2, forming an important pillar of future research under the abstract umbrella it provides.
@@
LLM agents for complex problem solving. The rapid increase in the capabilities of LLMs has triggered multiple recent efforts to integrate them at the core of autonomous agents that interact with the environment, plan, and act to solve complex problems (Wang et al., 2024). Typical approaches adopt integrating different tool-usage capabilities into LLMs via efficient prompting, often with multimodal capabilities (Chen et al., 2023). A single agent is often insufficient to solve complex problems; thus, multiple agents with different capabilities have to be integrated. Recent efforts in LLM-based multi-agent systems seek to mimic such cooperative problem solving by role-playing LLMs via in-context examples (Li et al., 2023b) or fine-tuning (Juneja et al., 2024). Despite some promising achievements, such agentic ecosystems are inherently limited by the constituent LLMs’ inability to plan and execute tasks (Xie et al., 2024; Kambhampati et al., 2024). These frameworks of LLM-based autonomous agents are largely designed for autonomous problem solving, not for human-AI co-construction as we envision. Recently, arguments in favor of strategically allocating tasks between humans and LLM-based agents to exploit their distinct strengths have been put forward (He et al., 2024), which aligns with our approach of leveraging the strengths of both humans and AI agents in co-construction. An alternative view of our framework is hence an extension of LLM-based multi-agent systems with a human agent as a key component, focusing on the co-construction of solutions.
@@
Persistent Solution Space for Iterative Construction. A fundamental component of our proposed framework is the explicit representation of the construction space for systematic solution search. Such a persistent memory can be useful for LLM agents. Sumers et al. (2024) propose a cognitive architecture for language agents that connects LLMs to internal memory and external environments, grounding them in existing knowledge or external observations. Similarly, Modarressi et al. (2024) introduce a structured memory component that LLM agents can use for storage and retrieval. In terms of deployed products, Anthropic’s artifacts [^5] https://support.anthropic.com/en/articles/9487310-what-are-artifacts-and-how-do-i-use-them add an explicit representation of an LLM-constructed artifact to the Claude series of language models, which can be iterated on through further interaction. Although these approaches do not directly address expert-AI co-construction challenges, they relate to our approach by providing agents with persistent memory to store intermediate solutions and relevant information for problem solving.
@@ 3 HAI-Co2: Human-AI co-construction through preference-based search
In this section, we propose HAI-Co2, a framework for cooperative problem solving. Broadly speaking, HAI-Co2 is meant to formalize an interactive problem solving scenario, in which a human expert seeks to (co-)construct a candidate solution – such as a computer program – with the help of an AI agent. The problem solving process is conceived as a process of systematic search in a space $\mathcal{X}$ of candidate solutions, i.e., as a co-constructive process, in which candidate solutions are modified or extended step by step until a (sufficiently) good solution has been found. Therefore, we also refer to the search space $\mathcal{X}$ as the construction space. The construction space, its hierarchical organization and its topology (or neighborhood structure) are depicted in Figure 2.
@@ 3.1 Construction space and abstraction hierarchy
The construction space will typically be large, most often even (countably) infinite. For example, the construction space may consist of all computer programs in a specific programming language. Spaces of this kind cannot be specified in an explicit way. Instead, they will be defined implicitly and may even be adapted or designed on-the-fly in the course of the problem solving process. In this regard, the formal representation of candidate solutions is of major importance and will strongly influence the efficacy and efficiency of the problem solving process. Moreover, it is also clear that the representation of solutions will not be universal but rather specific to the expert domain. For example, a computer program will not be represented in the same way as a machine learning pipeline or data science workflow. It should be noted that we do not make any assumption of completeness for candidate solutions: at any stage of the search, a candidate solution ${\bm{x}}\in{\mathcal{X}}$ can be partial or incomplete (i.e., an incomplete codebase, an incomplete ML pipeline, etc.).
During problem solving, it is often useful to look at (candidate) solutions on multiple levels of abstraction. In many cases, for example, a rough draft of the solution is found in a first phase of the process, and this draft is then worked out in more detail in a second phase. More generally, one can imagine a search process that switches back and forth between different levels of abstraction whenever appropriate. Therefore, we assume the construction space ${\mathcal{X}}$ is equipped with a hierarchy of abstraction levels.
Formally, this can be modeled by a sequence ${\mathcal{X}}_{0},{\mathcal{X}}_{1},\ldots,{\mathcal{X}}_{J}$ of spaces, where ${\mathcal{X}}_{j}$ is a refinement of ${\mathcal{X}}_{j-1}$ – or, vice versa, ${\mathcal{X}}_{j-1}$ an abstraction of ${\mathcal{X}}_{j}$. We describe the abstraction process from $\mathcal{X}_{j}$ to $\mathcal{X}_{j-1}$ as a surjection $f_{j}:{\mathcal{X}}_{j}\rightarrow{\mathcal{X}}_{j-1}$ such that ${\bm{x}}^{\prime}=f_{j}({\bm{x}})\in{\mathcal{X}}_{j-1}$; that is, ${\bm{x}}^{\prime}$ is the abstraction of ${\bm{x}}$ on the abstraction level modeled by ${\mathcal{X}}_{j-1}$. We denote by $f^{(-1)}_{j}({\bm{x}}^{\prime})=\{{\bm{x}}\in{\mathcal{X}}_{j}\,|\,f_{j}({\bm{x}})={\bm{x}}^{\prime}\}$ the set of all refinements of ${\bm{x}}^{\prime}\in{\mathcal{X}}_{j-1}$ on abstraction level ${\mathcal{X}}_{j}$. Note that refinements are not unique, which is why a transition from ${\mathcal{X}}_{j-1}$ to ${\mathcal{X}}_{j}$ may come with a certain arbitrariness.
In our case study on code generation presented in Section 5, we implement the construction space on three levels of abstraction, considering a Python program as a refinement of a UML diagram, which in turn is a refinement of a natural language specification. The refinement maps are implemented by suitably prompted LLMs that take a representation in a higher abstraction space as input and produce a solution representation in the lower level of abstraction as output.